# Reconstructing the whole 6D properties of the Sagittarius stream with N-body simulations

### 19 Dec 2023 v2

@@ Abstract
It is a challenge to reproduce the full 6D space-phase properties of Sagittarius (Sgr) dwarf galaxy and its Stream simultaneously. Using N-body simulations with a Milky Way mass of 5.2 M and a “scaling down” Sgr mass of 9.3 M, from a qualitative point of view, we have been able to reproduce well all 3D spatial features of Sgr stream, including its core, leading and trailing arms, and their associated bifurcations, moreover, the overall trend of the reported 3D kinematics properties of the Sgr stream have also been reproduced without fine tuning. Furthermore, we also find that our model fails in reproducing the exact behaviours of the line-of-sight velocity and angular-energy distributions. It let us to suggest that significant further progress might be achievable after introducing a major component in the Sgr progenitor, which is the gas that dominates all Irregular dwarf galaxies in the Sgr mass range and can slow down the radial velocity of Sgr before its removal, if gas can not solve this problem then we will consider a non-spherical Milky Way halo with hot gas, LMC, etc. As the first step for us to towards the complete understanding of the Sgr system, this progress is also advancing our understanding of the bifurcations, the generation of which might be due to the MW shocks at each pericenter passage, and also be linked to the orientation and disk-shape in the initial conditions.
@@
The Sagittarius dwarf spheroidal galaxy111We have ultimately decided to defer the modeling of the full Sagittarius dwarf galaxy with gas to a later contribution, which will also be part of the China-France collaboration on the Local group. So please note that this paper II has been delayed to contribute in the future (Paper I: arXiv:2205.02306v3). is the third most massive system surrounding the Milky Way. It has been discovered by Ibata et al. (1994), and its associated stellar stream a few years later (Ibata et al., 2001; Yanny et al., 2000; Newberg et al., 2002; Majewski et al., 2003). Majewski et al. (2003) using the 2MASS dataset revealed the presence of two arms of the stream in the south and north sides. With the help of SDSS data-set, precession of the Sagittarius stream and different evolution behaviours of the primary (brighter) tails and the secondary (fainter) tails both in the North and the South were unveiled in Belokurov et al. (2014). The current position and velocity of the Sgr centre are [X, Y, Z] = [17.5, 2.5, 6.5] kpc, V = [237.9, 24.3, 209.0] km s and the Sgr distance from the Sun is around 26.5 kpc according to Vasiliev et al. (2020) based on observational data222X, Y, and Z are in the left-handed Galactocentric cartesian coordinate same as this work, the sun location is at the left side of the Galactic centre..
@@
Most models cannot reproduce the stream bifurcations (see Fig. 4 of Ramos et al. 2022), including the model of Law & Majewski (2010) that matches well the Gaia DR2 data, and the more recent model by Vasiliev et al. (2021). Vasiliev et al. (2021) suggested that the model with a massive LMC can match well misalignment between stream track and sky-plane velocity in the Sgr coordinate, suggesting us that it might require a LMC mass of about 1.3 M associated to a Milky Way mass near 5.6 M within 100 kpc and virial mass is about 9 M. The observed bifurcation is not fully retrieved and without discussing gas physics influence as mentioned in Vasiliev et al. (2021). Moreover, they used two models with/without LMC to reproduce the Sgr core and stream, perhaps it is reasonable to ask for us now whether one model without physical changes but considering the gas contribution can solve the misalignment problems and leading arm overestimation issues.
@@ 2 N-body simulation and model
Simulations in this paper are carried out by using GIZMO (Hopkins et al., 2015). N-body realisation of initial conditions are created using Zeno333https://home.ifa.hawaii.edu/users/barnes/software.html. In our N-body simulations, the Milky Way particle mass resolution is 1.375 M with 1.2 Million particles and Sgr mass resolution is 1.375 M with 18 Million particles.
@@
Table 1 details the initial conditions of our dynamical simulations. We note that the mass adopted for the Milky Way is consistent with its rotation curve as it is included in the Jiao et al. (2021, 2023) range, and is very close to that recently found by Wang et al. (2022), both of which mass range are confirmed indirectly by recent Gaia DR3 disk rotation curve work with the slope of 2.3 0.2 km s kpc (Wang et al., 2023a; Jiao et al., 2023). The adopted mass of the Milky Way here is considerably lower than most current estimates (Wang et al., 2020).
@@
| Particle mass (star/dark matter) | 1.375(5.575) | 1.375(5.513) | M |
| Dark Matter mass | 4.8 | 9.0 | M |
@@
| Stellar disc mass | 3.58 | 3.0 | M |
@@ 3 Observational Data
and are the longitude and latitude in the Sgr stream coordinate system very similar to Majewski et al. (2003). In this work, the leading arm is from 60 to 180, the trailing arm covers a range of 180 to 360, small differences for this will not change the conclusion here. For the Milky Way coordinates in this work, the core/stream is projected in a left-handed Galactocentric cartesian coordinates [X, Y, and Z]. We have adopted the solar motion from Drimmel & Poggio (2018) with [ ] = [12.9, 12.6, 7.78] km s. The circular speed of the Local Standard of Rest is adopted to be 233 km s (Drimmel & Poggio, 2018). The distance of the Sun from the Galactic center is chosen to be = 8.277 kpc (GRAVITY Collaboration et al., 2022) and = 20.8 pc (Bennett & Bovy, 2019), and we have verified that different solar motions and coordinates would not change our final conclusions, either.
@@ 4 Results
Figure 3 presents a detailed comparison of our model with observational data. It shows the projection of the stream in 3 planes (XY, YZ, XZ), X, Y, and Z are in the conventional Galactocentric cartesian coordinate. Top, middle, and bottom panels show the modeled stars, the observed RGB star distribution, and the overlap of the data and model, respectively. In all panels the core location is identified by the density peak value (red colour), similarly as in Vasiliev et al. (2021) for the location, and at about [X, Y, Z] = [17.5, 2.5, 6.5] kpc. The size is relatively smaller due to that we start from a “scaling down” Sgr progenitor so our velocity dispersion is a bit smaller by few km s than Vasiliev et al. (2020). It is can be noticed that the trailing arm in our model is more elongated than that shown by the Vasiliev et al. (2021) RGB stars. Moreover, it can be seen that we have reconstructed very clear north bifurcation at the top and bottom left panel, which is shown around X = 30 kpc, and some south bifurcation signal might be also shown around X = 60 kpc. Note that it starts to be bifurcated from around 20 to 30 kpc, which is not the same as previous works in the conmmunity, except the bifurcation sigals, we have no other better explanation for the moment.
@@ 5.1 Model comparisons
In our new model we don’t consider the LMC for the moment, which is not the main target at this stage, but perhaps we can attempt to propose a new possibility, that is to say, the misalignment might be caused by the gas (energy), which will be investigated more in our series of works. Other possibilities such like rotated milky way halo potential for the misalignment was also proposed recently (Valluri et al., 2021) so we think it is non-trival here to propose this new possibility that the misaligment might be solved by considering the cold gas of the Sgr, hot and cold gas of the Milky Way, time-dependent non-symmetrical potential in the future.
@@
Moreover, we find that both Milky Way and Sgr masses are relatively modest in our study when compared to that of Vasiliev et al. (2021) model. About the mass of the Milky Way, as known the Sgr initial masses of our model are smaller than that of Vasiliev et al. (2021) model with =2M and =36M. They claimed that half of the stars are stripped by the Milky Way, and also pointed out that initial mass of the Sgr could be even much higher to a few M according to cosmological abundance (Jiang & Binney, 2000). Besides this, the Law & Majewski (2010) model assumed an Sgr total initial mass of 6.4M, which is smaller than that in our modeling, but for which no dark matter halo has been assumed. According to Ibata et al. (2020), the Law & Majewski (2010) model matches well the Sagittarius stream properties. It seems that high initial masses for Sgr do not favor the reproduction of the stream, as also noticed by Ibata et al. (2020). We are aware that current work does not allow to give strong constraints on the Sgr initial mass, and we acknowledge that we have not fully investigated the whole mass ranges for both stars and dark matter.
@@ 5.2 Disk interactions
Finally, the interactions of the Sgr and the Milky Way will possibly excite, flare, warp, ridge, radial motions, bulk motions, oscillations, star formation and so on for the Galactic disc (Laporte et al., 2018, 2019, 2020; Bland-Hawthorn et al., 2019; Wang et al., 2018a, b, 2019, 2020a, 2020c, 2020d, 2023a; Yu et al., 2021; Antoja et al., 2021; Khoperskov & Gerhard, 2022; Recio-Blanco et al., 2022; Bennett et al., 2022; Yang et al., 2023; Li et al., 2023), and after a quick investigation during our simulation, we have been able to detect flaring features, which is encouraging us to work more about Sgr and Milky Way Disc topics including the chemo-dynamics.
@@ 6 Conclusions
Our modelling has adopted to relatively modest mass for the Milky Way total mass (5.2 M) that matches well its observational rotation curve (Jiao et al., 2021; Wang et al., 2023a, see references therein) robustly established using Gaia DR2 data (Eilers et al., 2019). The same applies for the “scaling down” Sgr total mass (9.3 M) with the halo scale radius is 1.6 kpc, which has been modeled as a disc system, together with a relatively massive halo. The variety of masses from different models ((Law & Majewski, 2010; Ibata et al., 2001; Vasiliev et al., 2021) and reference therein) indicates that we are far from being able to predict mass using the present modelling of the Sgr stream. For our modelling, if one would like to consider gas or adopt large scale radius of the dark matter halo, it could naturally be to get a bit higher initial mass, but the change will affect the MW components and initial orbits and infall time.

### 20 Mar 2026 v3

@@ Abstract
It is a challenge to reproduce the full 6D space-phase properties of Sagittarius (Sgr) dwarf galaxy and its Stream simultaneously. Using N-body simulations with a Milky Way mass of 5.2 M⊙ and a “scaling down” Sgr mass of 9.3 M⊙, from a qualitative point of view, we have been able to reproduce well all 3D spatial features of Sgr stream, including its core, leading and trailing arms, and their associated bifurcations, moreover, the overall trend of the reported 3D kinematics properties of the Sgr stream have also been reproduced without fine tuning. Furthermore, we also find that our model fails in reproducing the exact behaviours of the line-of-sight velocity and angular-energy distributions. It let us to suggest that significant further progress might be achievable after introducing a major component in the Sgr progenitor, which is the gas that dominates all Irregular dwarf galaxies in the Sgr mass range and can slow down the radial velocity of Sgr before its removal, if gas can not solve this problem then we will consider a non-spherical Milky Way halo with hot gas, LMC, etc. As the first step for us to towards the complete understanding of the Sgr system, this progress is also advancing our understanding of the bifurcations, the generation of which might be due to the MW shocks at each pericenter passage, and also be linked to the orientation and disk-shape in the initial conditions.
@@
The Sagittarius dwarf spheroidal galaxy111We have ultimately decided to defer the modeling of the full Sagittarius dwarf galaxy with gas to a later contribution, which will also be part of the China-France collaboration on the Local group. So please note that this paper II has been delayed to contribute in the future (Paper I: arXiv:2205.02306v3). is the third most massive system surrounding the Milky Way. It has been discovered by Ibata et al. (1994), and its associated stellar stream a few years later (Ibata et al., 2001; Yanny et al., 2000; Newberg et al., 2002; Majewski et al., 2003; Wang et al., 2024). Majewski et al. (2003) using the 2MASS dataset revealed the presence of two arms of the stream in the south and north sides. With the help of SDSS data-set, precession of the Sagittarius stream and different evolution behaviours of the primary (brighter) tails and the secondary (fainter) tails both in the North and the South were unveiled in Belokurov et al. (2014). The current position and velocity of the Sgr centre are [X, Y, Z] = [17.5, 2.5, 6.5] kpc, VX,Y,Z = [237.9, 24.3, 209.0] km s-1 and the Sgr distance from the Sun is around 26.5 kpc according to Vasiliev et al. (2020) based on observational data222X, Y, and Z are in the left-handed Galactocentric cartesian coordinate same as this work, the sun location is at the left side of the Galactic centre..
@@
Most models cannot reproduce the stream bifurcations (see Fig. 4 of Ramos et al. 2022), including the model of Law & Majewski (2010) that matches well the Gaia DR2 data, and the more recent model by Vasiliev et al. (2021). Vasiliev et al. (2021) suggested that the model with a massive LMC can match well misalignment between stream track and sky-plane velocity in the Sgr coordinate, suggesting us that it might require a LMC mass of about 1.3 M⊙ associated to a Milky Way mass near 5.6 M⊙ within 100 kpc and virial mass is about 9 M⊙. The observed bifurcation is not fully retrieved and without discussing gas physics influence as mentioned in Vasiliev et al. (2021). Moreover, they used two models with/without LMC to reproduce the Sgr core and stream, perhaps it is reasonable to ask for us now whether one model without physical changes but considering the gas contribution can solve the misalignment problems and leading arm overestimation issues.
@@ 2 N-body simulation and model
Simulations in this paper are carried out by using GIZMO (Hopkins et al., 2015). N-body realisation of initial conditions are created using Zeno333https://home.ifa.hawaii.edu/users/barnes/software.html. In our N-body simulations, the Milky Way particle mass resolution is 1.375 M⊙ with 1.2 Million particles and Sgr mass resolution is 1.375 M⊙ with 18 Million particles.
@@
Table 1 details the initial conditions of our dynamical simulations. We note that the mass adopted for the Milky Way is consistent with its rotation curve as it is included in the Jiao et al. (2021, 2023) range, and is very close to that recently found by Wang et al. (2022), both of which mass range are confirmed indirectly by recent Gaia DR3 disk rotation curve work with the slope of 2.3 0.2 km s-1 kpc-1 (Wang et al., 2023a; Jiao et al., 2023). The adopted mass of the Milky Way here is considerably lower than most current estimates (Wang et al., 2020).
@@
| Particle mass (star/dark matter) | 1.375(5.575) | 1.375(5.513) | M⊙ |
| Dark Matter mass | 4.8 | 9.0 | M⊙ |
@@
| Stellar disc mass | 3.58 | 3.0 | M⊙ |
@@ 3 Observational Data
and are the longitude and latitude in the Sgr stream coordinate system very similar to Majewski et al. (2003). In this work, the leading arm is from 60∘ to 180∘, the trailing arm covers a range of 180∘ to 360∘, small differences for this will not change the conclusion here. For the Milky Way coordinates in this work, the core/stream is projected in a left-handed Galactocentric cartesian coordinates [X, Y, and Z]. We have adopted the solar motion from Drimmel & Poggio (2018) with [ ] = [12.9, 12.6, 7.78] km s-1. The circular speed of the Local Standard of Rest is adopted to be 233 km s-1 (Drimmel & Poggio, 2018). The distance of the Sun from the Galactic center is chosen to be = 8.277 kpc (GRAVITY Collaboration et al., 2022) and = 20.8 pc (Bennett & Bovy, 2019), and we have verified that different solar motions and coordinates would not change our final conclusions, either.
@@ 4 Results
Figure 3 presents a detailed comparison of our model with observational data. It shows the projection of the stream in 3 planes (XY, YZ, XZ), X, Y, and Z are in the conventional Galactocentric cartesian coordinate. Top, middle, and bottom panels show the modeled stars, the observed RGB star distribution, and the overlap of the data and model, respectively. In all panels the core location is identified by the density peak value (red colour), similarly as in Vasiliev et al. (2021) for the location, and at about [X, Y, Z] = [17.5, 2.5, 6.5] kpc. The size is relatively smaller due to that we start from a “scaling down” Sgr progenitor so our velocity dispersion is a bit smaller by few km s-1 than Vasiliev et al. (2020). It is can be noticed that the trailing arm in our model is more elongated than that shown by the Vasiliev et al. (2021) RGB stars. Moreover, it can be seen that we have reconstructed very clear north bifurcation at the top and bottom left panel, which is shown around X = 30 kpc, and some south bifurcation signal might be also shown around X = 60 kpc. Note that it starts to be bifurcated from around 20 to 30 kpc, which is not the same as previous works in the conmmunity, except the bifurcation sigals, we have no other better explanation for the moment.
@@ 5.1 Model comparisons
In our new model we don’t consider the LMC for the moment, which is not the main target at this stage, but perhaps we can attempt to propose a new possibility, that is to say, the misalignment might be caused by the gas (energy), which will be investigated more in our series of works. Other possibilities such like rotated milky way halo potential for the misalignment was also proposed recently (Valluri et al., 2021) so we think it is non-trival here to propose this new possibility that the misaligment might be solved by considering the cold gas of the Sgr, hot and cold gas of the Milky Way (Kalberla & Kerp, 2009; Barnes, 2002), time-dependent non-symmetrical potential in the future.
@@
Moreover, we find that both Milky Way and Sgr masses are relatively modest in our study when compared to that of Vasiliev et al. (2021) model. About the mass of the Milky Way, as known the Sgr initial masses of our model are smaller than that of Vasiliev et al. (2021) model with =2M⊙ and =36M⊙. They claimed that half of the stars are stripped by the Milky Way, and also pointed out that initial mass of the Sgr could be even much higher to a few M⊙ according to cosmological abundance (Jiang & Binney, 2000). Besides this, the Law & Majewski (2010) model assumed an Sgr total initial mass of 6.4M⊙, which is smaller than that in our modeling, but for which no dark matter halo has been assumed. According to Ibata et al. (2020), the Law & Majewski (2010) model matches well the Sagittarius stream properties. It seems that high initial masses for Sgr do not favor the reproduction of the stream, as also noticed by Ibata et al. (2020). We are aware that current work does not allow to give strong constraints on the Sgr initial mass, and we acknowledge that we have not fully investigated the whole mass ranges for both stars and dark matter.
@@ 5.2 Disk interactions
Finally, the interactions of the Sgr and the Milky Way will possibly excite, flare, warp, ridge, radial motions, bulk motions, oscillations, star formation and so on for the Galactic disc (Laporte et al., 2018, 2019, 2020; Bland-Hawthorn et al., 2019; Wang et al., 2018a, b, 2019, 2020a, 2020c, 2020d, 2023a; Yu et al., 2021; Antoja et al., 2021; Khoperskov & Gerhard, 2022; Recio-Blanco et al., 2022; Bennett et al., 2022; Yang et al., 2023; Li et al., 2023; Wang et al., 2019), and after a quick investigation during our simulation, we have been able to detect flaring features, which is encouraging us to work more about Sgr and Milky Way Disc topics including the chemo-dynamics.
@@ 6 Conclusions
Our modelling has adopted to relatively modest mass for the Milky Way total mass (5.2 M⊙) that matches well its observational rotation curve (Jiao et al., 2021; Wang et al., 2023a, see references therein) robustly established using Gaia DR2 data (Eilers et al., 2019). The same applies for the “scaling down” Sgr total mass (9.3 M⊙) with the halo scale radius is 1.6 kpc, which has been modeled as a disc system, together with a relatively massive halo. The variety of masses from different models ((Law & Majewski, 2010; Ibata et al., 2001; Vasiliev et al., 2021) and reference therein) indicates that we are far from being able to predict mass using the present modelling of the Sgr stream. For our modelling, if one would like to consider gas or adopt large scale radius of the dark matter halo, it could naturally be to get a bit higher initial mass, but the change will affect the MW components and initial orbits and infall time.